### v2

@@ Abstract
Mixture-of-Experts (MoE) has emerged as a practical approach to scale up parameters for the Transformer model to achieve better generalization while maintaining a sub-linear increase in computation overhead. Current MoE models are mainly built with expert parallelism on distributed devices. However, it usually depends on homogeneous devices to deploy and suffers from heavy communication overhead and computation redundancy. In this paper, we explore developing a Heterogeneous-aware EXpert Allocation framework, HEXA-MoE, with significantly enhanced computing efficiency. It contains two components: ($1$) Expert-Specific Operators. We replace the typical general matrix multiplication or grouped matrix multiplication interfaces with our operators, which allows the computing to be performed in an in-place manner with ZERO redundancy. ($2$) Adaptive Data- and Model-Centric Configurations for different workload scales. Specifically, we introduce a pipeline-shared cache on each device to tackle the heavy memory consumption in the existing data-centric MoE library. Comprehensive experiments on the Swin-MoE benchmark consistently reveal the effectiveness of our HEXA-MoE framework, i.e., reducing $10\%\sim 48\%$ memory consumption and achieving 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to current state-of-the-art MoE libraries. Furthermore, we examine our HEXA-MoE with heterogeneous devices for both data- and model-centric settings. Promising results show that employing optimal parallel configuration with HEXA-MoE on heterogeneous devices can substantially minimize overall latency. Codes are available at here.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Transformers have become the de facto architecture for a vast range of machine learning tasks including natural language process Vaswani (2017); Raffel et al. (2020), computer vision Liu et al. (2021); He et al. (2022b) and multi-modal learning Li et al. (2022); Liu et al. (2024). To scale up model parameters for stronger learning capacity and better generalization following the scaling law Kaplan et al. (2020), Mixture-of-Experts (MoE) Jiang et al. (2024) has been demonstrated as a practical and computation-efficient approach. For a single MoE layer, the feed-forward network (FFN) is expanded to a group of experts compared to the dense counterpart, activated selectively during computing. This dynamic nature poses peculiar challenges to system design. Current MoE frameworks usually present some unique attributes that differentiate them from dense models:
@@
  1. Distributed computing is required for MoE due to its sheer size, and expert parallelism Lepikhin et al. (2020) is the most commonly used technique, which distributes the experts in one layer on different devices.
@@
  2. Since the workload varies dynamically for each expert in each layer, we have to re-organize the local data batches for expert computing via dispatch and combine operations. Specifically, we have to calibrate the workload for each expert to utilize the general matrix multiplication (GeMM) or grouped GeMM interface.
  3. Since each token may activate any one or several experts and different experts may be distributed on different devices in expert parallelism, it is necessary to communicate among devices before and after each MoE layer via all-to-all communication.
Figure 2: Tensor Parallelism vs Expert Parallelism for MoE layer. We take $4$ experts on $2$ devices as an example and visualize the parameter distribution of $2$ MLPs. For expert parallelism, each device can only keep several experts, while for tensor parallelism the parameter allocation is more flexible since the partition of parameters is along the hidden dimension.
@@
These peculiar designs lead to inefficiency for both training and inference in MoE computing. On the one hand, dispatch and combine operations result in unnecessary token padding or discarding Gale et al. (2023), leading to both redundant computation and compromise on model performance. Meanwhile, conventional expert parallelism depends on all-to-all communication, which requires synchronization and occupies a lot of latency Li et al. (2023). On the other hand, current expert parallelism for MoE models is mainly deployed on homogeneous devices. However, a cluster equipped with the latest homogeneous devices is usually much more expensive than heterogeneous devices due to the iteration of GPU development and the decrease on prices for old-version GPUs. If we want to adapt expert parallelism to heterogeneous devices, it would be necessary to rearrange the expert distribution to utilize the total computing capacity better. However, the inherent dynamic property of the MoE layer makes it difficult to implement, as the workload of each expert changes dynamically in each step during computing He et al. (2022a); Li et al. (2023). Meanwhile, since heterogeneous devices are cheaper and easier to access if MoE computing can be re-designed to be heterogeneous-aware, it would benefit wider deployment.
@@
This paper proposes HEXA-MoE, a heterogeneous-aware framework for efficient MoE computing. We first find that if compiling the MoE layer into GeMM like other operators such as Linear, Convolution, etc., the unnecessary components such as token padding or discarding are unavoidable. To tackle it, we investigate the forward and backward propagation for the MoE layer thoroughly, and propose to replace current GeMM or grouped GeMM Gale et al. (2023) interfaces with expert-specific operators. Specifically, all the computations related to the MoE layer can be summarized as built on 3 operators, i.e., expert-specific matrix multiplication (ESMM), expert-specific summation (ESS) and expert-specific transposed matrix multiplication (ESTMM). The forward propagation only requires ESMM, while backward requires all the $3$ operators. We implement them in a highly-parallelized manner on GPU, and distribute the MoE layer with tensor parallelism Narayanan et al. (2021) as an alternative to expert parallelism, shown in Figure 2.
@@
HEXA-MoE also considers different workload scales for MoE computing and is adapted to both data- and model-centric configurations. Specifically, the difference lies in the content of communication. For model-centric, local data batches are synchronized among devices, and model parameters are kept still, while for data-centric, local parameters are gathered to each device, while local data batches are kept still Liu et al. (2023). When the scale of data is larger than the model parameters, the data-centric setting would perform better. However, previous work preserves all the gathered MoE parameters on each device for backward propagation, leading to huge memory consumption. To achieve better memory efficiency while maintaining high computing speed, we propose to introduce a pipeline-shared cache on each device, which is updated dynamically during both forward and backward propagation, and the all-gather communication can be overlapped with non-MoE operations such as attention. Essentially, a data-centric setting can be viewed as data parallelism, which is easy to adapt to heterogeneous devices. A model-centric setting is employed when the size of model parameters is larger than the data workload. In this case, all gathered communication is applied to the data batches, and tensor parallelism is applied to each MoE layer. Each device processes all the data batches using the parameter chunk it preserves. This can also be employed in heterogeneous devices via adaptively allocating expert parameters, as illustrated in Figure 2.
@@ 1. $\star$
    We propose to refactor MoE computing with expert-specific operators, enabling it to be performed in an in-place manner without the requirement of dispatch and combine operations. In this way, ZERO computation redundancy would be introduced. We implement these operators in a highly parallelized manner on GPU.
@@ 2. $\star$
    We develop HEXA-MoE based on the expert-specific operators for both data- and model-centric settings, considering different scales of workload, and propose to employ pipeline-shared cache for data-centric setting to improve memory efficiency. Experiments with Swin-MoE show that HEXA-MoE can reduce $10\%\sim 48\%$ memory usage and achieve 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to state-of-the-art MoE computing libraries.

### v4

@@ Abstract
Mixture-of-experts (MoE) has emerged as a practical approach to scale up parameters for Transformer model to achieve better generalization while keeping computational efficiency. Current MoE models are mainly deployed with expert parallelism on distributed devices, which typically depends on homogeneous devices, and suffers from heavy communication overhead as well as computation redundancy under scaled workload. To tackle these teasers, we propose a Heterogeneous-aware EXpert Allocation framework, HEXA-MoE, with significantly enhanced efficiency. It contains two components: ($1$) Expert-Specific Operators. We replace the typical GeMM or grouped GeMM interfaces with our proposed expert-specific operators, making expert computing to be performed in an in-place manner with almost no redundant FLOPs. ($2$) Adaptive Data- and Model-Centric configurations for different workload scales. We introduce a memory-efficient computation-communication overlapping scheme to tackle the heavy memory consumption in current data-centric libraries, which can accelerate training with heavy workloads. Comprehensive experiments on the Swin-MoE benchmark consistently reveal the effectiveness of HEXA-MoE, i.e., reducing $10\%\sim 48\%$ memory consumption and achieving 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to current state-of-the-art MoE libraries. We further examine HEXA-MoE on heterogeneous devices, and promising results show that employing optimal parallel configuration can better utilize global computation resources, and substantially minimize overall latency. Codes are available at here.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Convergence Analysis. HEXA-MoE can significantly surpass Tutel on MoE training due to the specialized designs.
@@
Transformers have become the de facto architecture for a vast range of machine learning tasks including natural language process (Vaswani, 2017; Raffel et al., 2020), computer vision (Liu et al., 2021; He et al., 2022b) and multi-modal learning (Li et al., 2022; Liu et al., 2024). To scale up model parameters for stronger learning capacity and better generalization following the empirical scaling law (Kaplan et al., 2020), Mixture-of-experts (MoE) (Jiang et al., 2024) has been demonstrated as a practical and computation-efficient approach. For a single Transformer layer, MoE expands the feed-forward network (FFN) to a group of experts with the same architecture and different parameters, activated selectively during computing. This dynamic nature poses peculiar challenges to system design. Current MoE frameworks usually present some unique attributes that differentiate them from dense models:
@@
  1. Distributed computing is required due to its sheer size, and expert parallelism (Lepikhin et al., 2020) is the most commonly used technique, which distributes the experts in one layer on different devices.
@@
  2. The dynamic workload for each expert makes it necessary to employ additional dispatch and combine operations in expert parallel to utilize the general matrix multiplication (GeMM) or grouped GeMM interface.
  3. Dispatch and combine rely on synchronous all-to-all communication to assign tokens to distributed experts.
@@
These peculiar designs lead to apparent inefficiency for MoE training. On the one hand, current MoE libraries are either static like Tutel (Hwang et al., 2023), or dynamic like MegaBlocks (Gale et al., 2023). The former depends on an additional hyper-parameter named expert capacity to calibrate the workload for different experts, which suffers from redundant memory allocation or access, while the latter tends to raise out of memory (OOM) error during runtime due to its uncertainty. Meanwhile, the synchronous all-to-all communication can occupy over $40\%$ runtime with scaled model and devices (Hwang et al., 2023). On the other hand, expert parallelism is mainly deployed on homogeneous devices. However, a homogeneous cluster with cutting-edge devices is much more expensive than heterogeneous ones due to the fast iteration of modern GPU and the decrease on prices for outdated ones. Adapting expert parallelism to heterogeneous devices would rely on rescheduling expert placement to utilize global computing resources sufficiently. However, the inherent dynamic property of MoE makes it difficult to implement, as the workload of each expert changes dynamically in each step (He et al., 2022a; Li et al., 2023). Since heterogeneous devices are cheaper and easier to access, if MoE training can be refactored to be heterogeneous-aware, it would advance wider deployment.
@@
We propose HEXA-MoE, a completely-static MoE training framework with minimized memory consumption and heterogeneous-awareness. We first find that if implementing MoE layer with GeMM, the unnecessary components such as token padding or discarding are unavoidable, while grouped GeMM (Gale et al., 2023) would lead to dynamic runtime behavior and uncertainty. To tackle it, we investigate the forward and backward propagation for the MoE layer thoroughly, and propose to replace (grouped) GeMM with expert-specific operators. Specifically, the basic MoE operations are essentially build upon $3$ operators, i.e., expert-specific matrix multiplication (ESMM), summation (ESS) and transposed matrix multiplication (ESTMM). The forward propagation only requires ESMM, while backward requires all. We implement them with specialized GPU kernels, and distribute the MoE layer with tensor parallelism (Narayanan et al., 2021) as an alternative to expert parallelism, which is easy to be heterogeneous-aware.
@@
HEXA-MoE also considers different workload scales for MoE training, and is adapted to both data- and model-centric configurations. Their difference lies in the content of communication. For model-centric, local mini-batches are synchronized among devices and model parameters are kept still, while for data-centric, local parameters are gathered for each device, while local mini-batches are kept still (Liu et al., 2023). Data-centric setting outperforms model-centric with scaled workloads. However, previous library preserves all the gathered parameter shards on each device for backward pass, leading to huge memory consumption. To tackle this teaser, we design a novel memory-efficient communication-communication overlapping scheme utilizing a pipeline-shared cache on each device, which is updated dynamically for both forward and backward pass. Both settings are completely static in runtime, and the workload of each device can be easily determined from batch size or tensor parallelism configuration (sub-dimension of FFN intermediate size), making it easy to be adapted to heterogeneous devices via specialized expert allocation.
@@ 1. $\star$
    A completely static MoE training framework with minimized memory footprint. To our best knowledge, HEXA-MoE is the first MoE library with completely static and minimal memory footprint, enabling in-place computing with specialized kernel, and employing tensor parallelism for distributed training, delivering exact result with faster speed, as shown in Figure 1.
@@ 2. $\star$
    Advanced efficiency. HEXA-MoE is built upon our proposed expert-specific operators, which tackles scaled workload with data-centric strategy via memory-efficient communication-computation overlapping. Experiments on Swin-MoE show that HEXA-MoE can reduce $10\%\sim 48\%$ memory usage and achieve 0.54.3×0.5$\sim$ 4.3$\times$0.5 ∼ 4.3 × speed up compared to SOTA MoE libraries.